OpenStudy (anonymous):

@IMStuck Patti’s monthly pay is $3460.96. She usually uses 3/9 of that amount to pay her household expenses, and 1/4  of that amount to pay for tuition. She deposits the rest into her savings account. This month, Patti deposited twice the amount of money into her savings account than she normally deposits.   How much did she deposit into her savings account this month?

OpenStudy (imstuck):

Let's find out how much she usually deposits into savings by figuring out how much 3/9 and 1/4 is. Like this:

OpenStudy (anonymous):

7/12

OpenStudy (imstuck):

$3460.96 times 3/9 = 1153.65 That goes into household expenses. Then she puts 1/4 away for school, so 1/4 of $3460.96 is 865.24. Add those together and it comes to 2018.89. That leaves $3460.96-2018.89=1442.07

OpenStudy (imstuck):

$1442.07 is what she usually has left to put away for savings. It's telling you that this month she deposited twice the amount into her savings than normal. So how much did she deposit this month?

OpenStudy (anonymous):

$3460.96 times 3/9 = 1153.65 That goes into household expenses. Then she puts 1/4 away for school, so 1/4 of $3460.96 is 865.24. Add those together and it comes to 2018.89. That leaves $3460.96-2018.89=1442.07

OpenStudy (anonymous):

$1442.07 is what she usually has left to put away for savings. It's telling you that this month she deposited twice the amount into her savings than normal. So how much did she deposit this month?

OpenStudy (imstuck):

i see that you have a fraction for an answer? I don't know if it is correct, I didn't check it, but I think they want a dollar amount as an answer.

OpenStudy (anonymous):

1442.07 is the answer

OpenStudy (imstuck):

No it's not the answer. That 1442.07 is the amount she puts away for one month normally. They asked you how much she deposited if she put away twice that normal amount this month.

OpenStudy (anonymous):

No it's not the answer. That 1442.07 is the amount she puts away for one month normally. They asked you how much she deposited if she put away twice that normal amount this month.

OpenStudy (imstuck):

TY for the medal to whomever gave it to me. And i think that -StudyShark- is mimicking me?

OpenStudy (anonymous):

i gave u the medal

OpenStudy (imstuck):

TY

OpenStudy (anonymous):

and he is mimicking you lol

OpenStudy (imstuck):

I don't appreciate it. I'm too old for games like that.